A wheel is spinning at 5.38 rad/s, and has an angular momentum of 4.05 kg.m2/s. What is its moment of inertia?


1
Expert's answer
2021-02-12T16:17:08-0500

From the definition of angular momentum, we have:


"L=I\omega,""I=\dfrac{L}{\omega}=\dfrac{4.05\ \dfrac{kg\cdot m^2}{s}}{5.38\ \dfrac{rad}{s}}=0.75\ kg\cdot m^2."